(I don't think I understand what this equation is saying)

Conductivity = A*(1-exp(-c*x))+B

A= 2.385E+04

C= 3.685

B= -3.693

Find the conductivity at x= 0.1
Answer #1

This equation gives conductivity as a function of x

A,B,C and x values are given. Substituting them in the equation gives conductivity at that x.

Conductivity = 2.385*10^4 *(1 - exp(-3.685*0.1)) - 3.693

Conductivity = 2.385*10^4 * (1- 0.692) - 3.693

Conductivity = 7345.8 - 3.693 = 7.342*10^3